OpenStudy (geometry_hater):

The measure of an interior angle of a regular polygon is 140. Find the number of sides the polygon has.

OpenStudy (anonymous):

\[\frac{180(n-2)}{n} = 140\]

OpenStudy (anonymous):

180 n - 360 = 140n 40n = 360 n = 9 so it is 9 sided

OpenStudy (geometry_hater):

thanks